### v2

@@ Abstract
A growing number of central authorities use assignment mechanisms to allocate students to schools in a way that reflects student preferences and school priorities. However, most real-world mechanisms incentivize students to strategically misreport their preferences. In this paper, we provide an approach for identifying the causal effects of school assignment on future outcomes that accounts for strategic misreporting. Misreporting may invalidate existing point-identification approaches, and we derive sharp bounds for causal estimates that robustly account for strategic behavior. Our approach applies to any mechanism as long as there exist placement scores and cutoffs that characterize that mechanism’s allocation rule. We use data from a deferred acceptance mechanism that assigns students to more than 1,000 university–major combinations in Chile. Matching theory predicts that students’ behavior in Chile should be strategic because they can list only up to eight options, and we find empirical evidence consistent with such behavior. Our bounds are informative enough to reveal significant heterogeneity in graduation success with respect to preferences and school assignment.
@@ 1 Introduction
One of the most important decisions students make is their choice of field and institution of education. Identification of the impact of such choices on future outcomes is a critical step in the study of this decision process and of public education policies. The causal effects of schooling may vary widely across economic agents because of heterogeneous skills and preferences. Moreover, individuals’ expectations about potential returns may prompt them to choose schools strategically. Heterogeneity and selection make identification of causal effects very challenging, especially when students face a large number of unordered options. On the positive side, a growing number of schools use centralized assignment mechanisms, which offer credible instruments based on the discontinuities generated by random assignment of comparable students to different schools (kirkeboen2016 and abdul_ecta2022). Such discontinuities arise from the unpredictable admission cutoffs at which students are matched to schools under the assignment mechanisms.
@@
Many centralized school assignment mechanisms effectively amount to a quasi-experimental design where two groups of individuals who share similar scores are assigned to different schools based on how their scores relate to admission cutoffs. The assignment in a matching characterized by such a cutoff depends on the student’s preferences over feasible schools. Unlike in a typical regression discontinuity (RD) design, in this setting, students on the same side of the cutoff do not necessarily receive the same assignment. For example, individuals with similar scores just above a certain cutoff may all prefer to go to the same school $a$ but could have very different second-best options if they fall on the other side of the cutoff. In such a context, kirkeboen2016 construct comparable groups of individuals near a cutoff by conditioning on local preferences—that is, by selecting students whose preferences yield identical first- and second-best options if they fall, respectively, above and below that cutoff. Controlling for local preferences that equal an identical pair of schools, e.g., $(a,b)$, allows the RD to identify the causal effect of a change in the school assignment from $b$ to $a$, averaged over individuals who prefer $a$ over $b$.Although powerful, this design relies on the assumption that students report their true preferences to the centralized mechanism. A natural question to ask, then, is how often students submit their true preferences. In fact, most real-world school assignment mechanisms create incentives for students to misreport their preferences. agarwal2018 and fack2019 provide thorough discussions with several real-world examples of this.
@@
This paper derives sharp bounds for estimates of the causal effects of school assignment on future outcomes in mechanisms characterized by a cutoff and strategic student behavior. We propose a two-step identification approach that is robust to strategic reporting of preferences. In the first step, the researcher partially identifies local preferences and constructs local preference sets for each student. We provide several tools for constructing these sets in the context of a student-proposing deferred acceptance (DA) mechanism where constraints on the submitted preferences lead to strategic behavior. Outside such a context, researchers may employ alternative tools to partially identify local preferences, as the second step of our procedure does not require a particular method to be used in the first step. For example, the identification method of agarwal2018 applies to a general class of mechanisms with cutoff representation that includes variants of the DA, Boston mechanism, first preferences first, Chinese parallel mechanism, etc. Finally, in the second step, the researcher employs an RD identification strategy that controls for local preference sets and partially identifies the causal effects of school assignment.
@@
Strategic behavior among students depends on the characteristics of the assignment mechanism. A mechanism is said to be strategy proof if submitting true preferences is a weakly dominant strategy for all students. For example, dubins1981 demonstrate that the DA mechanism is strategy proof. However, this result breaks down when the mechanism imposes constraints on the preferences that students can submit. In many real-world school assignment mechanisms, the number of schools is too large for students to feasibly rank all schools. The central authorities running these systems may either limit the number of schools that students may rank or impose costs on the basis of the number of schools submitted. See Table 1 Panel B by fack2019 for examples.
@@
Our first-step tools for partial identification of local preferences naturally require assumptions on students’ strategic behavior. We motivate our assumptions following the important contributions of haeringer2009. They study a game where students submit constrained preference rankings and a central mechanism allocates the students to schools. One of their important findings is that it is rational for students to submit partial orders of their true preferences in some mechanisms. Specifically, suppose that a mechanism is strategy proof when students are free to rank any number of schools, as under, e.g., the unconstrained DA or top trading cycles (TTC) mechanisms. Then, if the preference rankings are constrained to having at most $K$ schools, a student can do no better than selecting $K$ schools among her acceptable schools and ranking them according to her true preferences.
@@
The key assumption for the first step in our method is that students submit only partial orders of their preferences. This feature is in addition to the cutoff characterization of the matching, which we assume throughout the paper. Cutoff characterization means that a student is matched to her best feasible school, where “best” is defined according to her true preferences and a school is feasible if the student’s placement score clears that school’s admission cutoff. Our assumption about this cutoff characterization is satisfied when the matching outcome is stable (azevedo2016).Stability means that each student is matched to an acceptable school and all slots in better schools have been filled with people who have better placement scores. In constrained DA and TTC mechanisms, stability occurs in Nash equilibrium of the preference revelation game under appropriate conditions on the placement scores (Theorems 6.3 and 6.4, haeringer2009). Stability occurs in Nash equilibrium without restrictions on the scores in the constrained serial dictatorship (SD) mechanism, which is a particular case of DA. We characterize sharp local preference sets for every individual that are compatible with the observed data and these model assumptions. Our random local preference sets contain the true local preference random variable with probability one. We also show a way to shrink these sets by imposing assumptions on students’ expectations regarding the outcome of the match and by assuming students maximize expected utility, in line with agarwal2018.
The form of our local preference sets has two important implications for applied work. First, if we ignore strategic behavior and condition only on reported local preferences, the RD strategy may yield inconsistent estimates. The reason is that when we select individuals with the same (strategically) reported local preferences, our local preference sets change discontinuously at a cutoff. The problem is akin to the manipulation problem in RD: we are controlling for a variable that is manipulable, that is, reported preferences. Second, suppose that we select individuals with reported local preferences $(a,b)$ and see that the local preference sets are all singletons $\{(a,b)\}$; in other words, our selection contains only individuals who do not manipulate. Even in this case, an RD strategy that controls for reported local preferences may still be inconsistent because the local preference sets may well contain the pair $(a,b)$ for someone whose reported local preferences differ from $(a,b)$. Thus, selecting individuals who report $(a,b)$ does not necessarily imply that we select all individuals whose true local preferences equal $(a,b)$.
@@
The second step of our approach relies on the local preference sets constructed in the first step with either our method or an alternative method. Given interest in a pair of schools $(a,b)$, we select all individuals whose local preference sets contain $(a,b)$ and whose placement scores are close to the cutoff for admission at school $a$. This subpopulation of individuals contains all individuals whose true local preferences equal $(a,b)$, but also other individuals. The average outcome in the subpopulation equals a weighted average of two averages: first, the average outcome for individuals with true local preferences $(a,b)$, which is interesting for the identification of causal effects, and second, the average outcome for individuals with true local preferences that differ from $(a,b)$. We do not know which individuals have preferences $(a,b)$, but we do characterize sharp bounds on the proportion of such individuals in the subpopulation using the random sets constructed in the first step. Thus, our setting fits the identification problem with corrupted data studied by horowitz1995. This method allows us to derive closed-form bounds on the first out of the two average outcomes above, which then leads to bounds on the average causal effects. This closed-form approach offers some intuition on when we can expect the bounds to be informative about or equal to the actual estimates (i.e., point identification). Although practical and intuitive, these closed-form bounds may not be sharp. Thus, building on molinari2020 and using random set theory, we characterize sharp bounds that are numerically computable when outcomes take finitely many values.
@@
Methods combining RD identification with school matching data have been popular among applied and theoretical researchers in economics (kirabo_ej2010; bertanha_joe2020). To the best of our knowledge, our paper is the first to prove RD identification of returns of school assignment in matching mechanisms with strategically reported preferences. We note that our two-step approach differs from the usual control function approach because our first step partially identifies the control variable instead of point-identifying it as in the usual approach. This paper unifies and complements two branches of the literature. One branch features the methods proposed by agarwal2018 and fack2019 that take into account strategic reporting and identify students’ true preferences, but they do not focus on causal effects of school matches on future outcomes. The other branch features the methods of kirkeboen2016 and abdul_ecta2022 that identify the causal effects of different assignments but control for reported instead of true preferences. chen2023nonparametric discusses methods applicable in the case when assignment is based both on lottery- and RD-driven variation.
@@
We apply our two-step identification strategy to matching data from Chile. Chile has a centralized DA mechanism that assigns students to university–major pairs. In 2010, 88,000 students ranked at most eight university–major pairs out of a total of 1,092 options available. Thus, the mechanism constrains students’ preference rankings, and students have incentives to behave strategically. The methods proposed by kirkeboen2016 and abdul_ecta2022 are hence not directly applicable in the Chilean case, even if all students report their preferences truthfully.
@@
The methods of kirkeboen2016 apply to the SD mechanism—a particular case of the DA mechanism in which all schools utilize the same placement score. In the SD case, the counterfactual set of schools for students just above and just below the cutoff does not vary across students. For example, for students just above the cutoff, their counterfactual set includes all schools whose cutoffs are lower than the cutoff in question. The same does not apply in DA, the mechanism used in Chile. Students have multiple placement scores, and the set of feasible schools may vary widely across students. Defining the counterfactual set is an important step in the RD identification strategy because local preferences are defined over these sets, which then become the control variable in the RD strategy.
@@
abdul_ecta2022 propose a solution to the DA counterfactual problem by constructing a propensity score control variable. They study New York City public high schools, where placement scores are functions of integer priority scores plus a continuously distributed variable with full support. This particular structure of school priorities does not correspond to the Chilean case, where program-specific placement scores are computed as functions of five primitive scores: math, language, history, science, and an average score from high school. These functions are different across programs and sometimes nonlinear. This setting requires the counterfactual set of schools to be carefully defined such that controlling for local preferences does not violate the continuity assumptions required by RD. We therefore propose a general method that applies to such empirical contexts and leads to point identification if students are truthful but partial identification otherwise.
We present several pieces of evidence that students in Chile behave strategically. In settings where behavior is strategic, existing methods that control for reported preferences are invalid, and identification requires the tools that we introduce in this paper. Our data contain more than three hundred university–major pairs for which the localized RD samples are large enough for our procedure to be implemented. For each pair, we construct bounds on the average effect of a change in the initial assignment between the second- and first-best options in that pair. We examine the effects on two binary outcomes: graduation from the first-best university–major and graduation from a top university. The bounds that we derive are informative about the sign of average treatment effects in 33–67% of the university–major pairs, depending on the outcome. The bounds are also informative in that they reveal substantial heterogeneity in the treatment effects across local preferences and second-best options. Finally, we compare our bounds to point estimates that we obtain by naively ignoring strategic behavior and simply controlling for reported preferences. The naive estimates fall outside our bounds for 25–57% of the university–major pairs, consistent with the fact that naive estimates are generally inconsistent for an average treatment effect parameter when students behave strategically.
@@
The rest of this paper proceeds as follows. Section 2 lays out the matching model for a continuum population of students and a finite number of schools. Section 3 examines point identification of average treatment effects when students are truth-tellers. Section 4 examines partial identification when students strategically report their preferences, with two subsections: Section 4.1 provides tools for construction of local preference sets that apply to constrained DA mechanisms, and Section 4.2 discusses how to use local preference sets constructed in this or other ways to derive bounds on the average treatment effects. We illustrate our identification approach with the Chilean data in Section 5. The appendix presents all proofs for the paper.
@@ 2 Model
We consider a continuum population of students and a set of $J$ schools, $\mathcal{J}:=\{1,\ldots,J\}$, that have capacities $\{q_{1},\ldots,q_{J}\}$ defined in terms of shares of the student population (azevedo2016). Denote by $\Omega$ the set of all students in the universe of interest, and use $\omega$ to index an individual student type. The student type consists of three objects. First, $Q(\omega)$ denotes the true (strict) preference relation of student $\omega$ over the set of options $\mathcal{J}^{0}:=\mathcal{J}\cup\{0\}$, which includes schools $\mathcal{J}$ and an outside option 00. For example, if $J=2$ and $Q(\omega)=\{1,2,0\}$, then 1 is preferred to 2 (i.e., $21Q(\omega)$ 2), 1 is preferred to 0 (i.e., $01Q(\omega)$ 0), and 2 is preferred to 0 (i.e., $02Q(\omega)$ 0). Let $\mathcal{Q}$ be the set of all strict preference relations over $\mathcal{J}^{0}$ that admit at least one school that is acceptable. A school $j\in\mathcal{J}$ is “acceptable” for student $\omega$ if it is preferred to that student’s outside option, i.e., $jQ(\omega)0$. We define $\bar{Q}$ as the weak preference relation induced by $Q$, i.e., $j\bar{Q}k\Leftrightarrow jQk$ or $j=k$. The second object of the student type is a vector of scores $\bm{\mathrm{R}}(\omega):=(R_{1}(\omega),\ldots,R_{J}(\omega))\in\mathcal{R}\subseteq\mathbb{R}^{J}$, where each school $j$ utilizes $R_{j}$ to rank students for admission. The third and last object, $Y(\omega,d)$, is the potential outcome of student $\omega$ if the student is assigned to option $d\in\mathcal{J}^{0}$. Each student has a potential outcome function $Y(\omega,\cdot)$ that maps from $\mathcal{J}^{0}$ to $\mathcal{Y}\subseteq\mathbb{R}$. We call $\bm{\mathrm{\Gamma}}$ the set of all possible potential outcome functions. The set of all student types is $\Omega:=\mathcal{Q}\times\mathcal{R}\times\bm{\mathrm{\Gamma}}$. In a continuum economy, there is a probability measure $\mathbb{P}$ over $\Omega$ and the Borel $\sigma$-algebra of the product space $\Omega$. We suppress the argument $\omega$ whenever it is unnecessary for ease of notation, e.g., $Y(d)$ vs. $Y(\omega,d)$ and $Q$ vs. $Q(\omega)$.
@@
A “matching” is described by a measurable function $\mu:\Omega\to\mathcal{J}^{0}$ that satisfies two conditions: for every $j\in\mathcal{J}$, (i) the mass of students matched to $j$ is less than or equal to the capacity of school $j$, i.e., $\mathbb{P}\{\omega:\mu(\omega)=j\}\leq q_{j}$; and (ii) the set of students who weakly prefer option $j\in\mathcal{J}^{0}$ over their matching, i.e., $\{\omega:j\bar{Q}(\omega)\mu(\omega)\}$, is an open set.444 azevedo2016 impose the same condition to rule out multiplicity of stable matchings that differ in a set of types with measure zero. For every student type $\omega$, $\mu(\omega)$ is either the school $j$ to which the student is matched or zero. When $\mu(\omega)=0$, the student is unmatched and takes an outside option. An important definition for this paper is that of stability.
@@
A mechanism $\varphi$ matches students to schools by mapping the students’ scores and submitted preference lists to schools. Student $\omega$ submits a preference list $P(\omega)\subseteq\mathcal{J}$, which is an ordered list of her acceptable schools. For example, for $J=3$, if $Q(\omega)=\{1,2,0,3\}$, then $P(\omega)=\{1,2\}$, as long as the student’s submitted list of acceptable schools reflects her true preferences. The number of schools in $P$, denoted $|P|$, is at least one because everyone participating in the match has at least one acceptable school. As with $\bar{Q}$, we also define $\bar{P}$ as the weak preference relation induced by $P$. A mechanism takes as inputs everyone’s submitted preferences (i.e., a correspondence $P:\Omega\rightrightarrows\mathcal{J}$) and everyone’s scores (i.e., a vector-valued function $\bm{\mathrm{R}}:\Omega\to\mathcal{R}$) and gives rise to a matching function. Formally, $\varphi(P,\bm{\mathrm{R}}):\Omega\to\mathcal{J}^{0}$. We say a mechanism $\varphi$ is strategy proof if submitting her true ranking of acceptable schools is weakly dominant for every student—in other words, if any student $\omega$’s misreporting of $P$ never leads to a better option and sometimes leads to a worse option, depending on what other students submit.We say a student is a truth-teller if her $P$ equals her true ranking of acceptable schools. Otherwise, we say she is strategic or not a truth-teller.
@@
For placement scores $\bm{\mathrm{S}}:\Omega\to\mathcal{S}\subseteq\mathbb{R}^{J}$, $\bm{\mathrm{S}}(\omega):=(S_{1}(\omega),\ldots,S_{J}(\omega))$ and admission cutoffs $\bm{\mathrm{c}}\in\mathcal{S}$, $\bm{\mathrm{c}}:=(c_{1},\ldots,c_{J})$, the set of feasible options of a student $\omega$ equals all schools for which her placement scores clear the admission cutoffs plus the outside option: $\{0\}\cup\{j\in\mathcal{J}:S_{j}(\omega)\geq c_{j}\}$; student $\omega$’s best feasible option is the option that ranks first according $Q(\omega)$ among her feasible options. We say the matching $\mu:\Omega\to\mathcal{J}^{0}$ has cutoff characterization if there exist placement scores $\bm{\mathrm{S}}:\Omega\to\mathcal{S}$ and admission cutoffs $\bm{\mathrm{c}}\in\mathcal{S}$ such that, for every $\omega\in\Omega$, the matching $\mu(\omega)$ equals student $\omega$’s best feasible option according to $Q(\omega)$.

### v3

@@ Abstract
A growing number of central authorities use assignment mechanisms to allocate students to schools in a way that reflects student preferences and school priorities. However, most real-world mechanisms incentivize students to strategically misreport their preferences. Misreporting complicates the identification of causal parameters that depend on true preferences, which are necessary inputs for a broad class of counterfactual analyses. In this paper, we provide an identification approach that is robust to strategic misreporting and derive sharp bounds on causal effects of school assignment on future outcomes. Our approach applies to any mechanism as long as there exist placement scores and cutoffs that characterize that mechanism’s allocation rule. We use data from a deferred acceptance mechanism that assigns students to more than 1,000 university–major combinations in Chile. Matching theory predicts and empirical evidence suggests that students behave strategically in Chile because they face constraints on their submission of preferences and have good a priori information on the schools they will have access to. Our bounds are informative enough to reveal significant heterogeneity in graduation success with respect to preferences and school assignment.
@@ 1 Introduction
One of the most important decisions students make is their choice of field and institution of education. Identification of the impact of such choices on future outcomes is a critical step in the study of this decision process and of public education policies. The causal effects of schooling may vary widely across economic agents because of heterogeneous skills and preferences. Moreover, individuals’ expectations about potential returns may prompt them to choose schools strategically. Heterogeneity and selection make identification of causal effects very challenging, especially when students face a large number of unordered options. On the positive side, a growing number of schools use centralized assignment mechanisms, which produce credible instruments based on the discontinuities generated by the assignment of comparable students to different schools (Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022)).
@@
Many centralized school assignment mechanisms effectively amount to a quasi-experimental design where two groups of individuals who share similar scores are assigned to different schools based on how their scores relate to admission cutoffs. The assignment in a matching characterized by such cutoffs depends on the student’s preferences over feasible schools. Unlike in a typical regression discontinuity (RD) design, in this setting, students on the same side of the cutoff do not necessarily receive the same assignment. For example, individuals with similar scores just above a certain cutoff may all prefer to go to the same “school jj” but could have very different second-best options if they fall on the other side of the cutoff. In such a context, Kirkeboen et al. (2016) construct comparable groups of individuals near a cutoff by conditioning on local preferences—that is, by selecting students whose preferences yield identical first- and second-best options if they fall, respectively, above and below that cutoff. Controlling for local preferences that equal a pair of schools, e.g., $(j,k)$, allows the RD to identify the causal effect of a change in the school assignment from $k$ to $j$, averaged over individuals who prefer $j$ over $k$.
Most real-world school assignment mechanisms create incentives for students to misreport their true preferences. Agarwal and Somaini (2018) and Fack et al. (2019) provide thorough discussions with several real-world examples of this. The empirical literature on school choice typically engages in counterfactual policy evaluations and an important class of such counterfactuals requires identification of causal parameters that are conditional on true preferences (Artemov et al. (2023), Section V.B). The identification of such causal parameters is the primary goal of our paper and we propose a novel RD strategy that controls for true local preferences. The challenge resides on the fact that true preferences are unobserved but may be partially identified under general forms of misreporting behavior.
@@
This paper derives sharp bounds for causal effects of school assignment on future outcomes in mechanisms with cutoff characterization and strategic student behavior. We devise a two-step identification approach that is robust to strategic reporting of preferences. In the first step, the researcher partially identifies local preferences and constructs local preference sets for each student. We provide several tools for constructing these sets in the context of a student-proposing deferred acceptance (DA) mechanism where constraints on the submitted preferences lead to strategic behavior. Outside such a context, researchers may employ alternative tools to partially identify local preferences, as the second step of our procedure does not require a particular method to be used in the first step. For example, the identification method of Agarwal and Somaini (2018) applies to a general class of mechanisms with cutoff representation that includes variants of the DA mechanism, Boston, First Preferences First, Chinese Parallel, etc. Finally, in the second step, the researcher employs an RD identification strategy that controls for local preference sets and partially identifies the causal effects of school assignment.
@@
Strategic behavior among students depends on the characteristics of the assignment mechanism. A mechanism is said to be strategy proof if submitting true preferences is a weakly dominant strategy for all students. For example, Dubins and Freedman (1981) demonstrate that the DA mechanism is strategy proof. However, this result breaks down when the mechanism imposes constraints on the preferences that students can submit. In many real-world school assignment mechanisms, the number of schools is too large for students to feasibly rank all schools. The central authorities running these systems may either limit the number of schools that students may rank or impose costs on the basis of the number of schools submitted (see Table 1 Panel B by Fack et al. (2019) for examples).
@@
Our first-step tools for partial identification of local preferences naturally require assumptions on students’ strategic behavior. We motivate our assumptions following the important contributions of Haeringer and Klijn (2009). They study a game where students submit constrained preference rankings and a central mechanism allocates the students to schools. One of their important findings is that it is rational for students to submit partial orders of their true preferences in some mechanisms. Specifically, suppose that a mechanism is strategy proof when students are free to rank any number of schools, as under, e.g., the unconstrained DA or Top Trading Cycles (TTC) mechanisms. Then, if the preference rankings are constrained to having at most $K$ schools, a student can do no better than selecting $K$ schools among her acceptable schools and ranking them according to her true preferences.
@@
The key assumption for our first-step tools is that students submit only partial orders of their preferences. This feature is in addition to the cutoff characterization of the matching, which we assume throughout the paper. Cutoff characterization means that a student is matched to her best feasible school, where “best” is defined according to her true preferences and a school is feasible if the student’s placement score clears that school’s admission cutoff. Our assumption about cutoff characterization is satisfied when the matching outcome is stable (Azevedo and Leshno, 2016). Stability means that each student is matched to an acceptable school and all slots in preferred schools have been filled with people who have better placement scores. In constrained DA and TTC mechanisms, stability occurs in Nash equilibrium of the preference revelation game under appropriate conditions on the placement scores (Theorems 6.3 and 6.4, Haeringer and Klijn (2009)). Stability occurs in Nash equilibrium without restrictions on the scores in the constrained serial dictatorship (SD) mechanism, which is a particular case of DA. We characterize sharp local preference sets for every individual that are compatible with the observed data and these model assumptions. Our random local preference sets contain the true local preference random variable with probability one. We also show a way to shrink these sets by imposing assumptions on students’ expectations regarding the outcome of the match and by assuming students maximize expected utility, in line with Agarwal and Somaini (2018).
@@
The second step of our approach relies on the local preference sets constructed in the first step with either our method or an alternative method. Given interest in a pair of schools $(j,k)$, we select all individuals whose local preference sets contain $(j,k)$ and whose placement scores are close to the cutoff for admission at school $j$. This subpopulation of individuals contains all individuals whose true local preferences equal $(j,k)$, but also other individuals. The average outcome in the subpopulation equals a weighted average of two averages: first, the average outcome for individuals with true local preferences $(j,k)$, which is interesting for the identification of causal effects, and second, the average outcome for individuals with true local preferences that differ from $(j,k)$. We do not know which individuals have preferences $(j,k)$, but we do characterize sharp bounds on the proportion of such individuals in the subpopulation using the random sets constructed in the first step. Thus, our setting fits the identification problem with corrupted data studied by Horowitz and Manski (1995). This method allows us to derive closed-form bounds on the first out of the two average outcomes above, which then leads to bounds on the average causal effects. This closed-form approach offers some intuition on when we can expect the bounds to be informative about or equal to the actual average causal effect (i.e., point identification). Although practical and intuitive, these closed-form bounds may not be sharp. Thus, building on Molinari (2020) and using random set theory, we characterize sharp bounds that are numerically computable when outcomes take finitely many values.
@@
Methods combining RD identification with school matching data have been popular among applied and theoretical researchers in economics (Jackson, 2010; Bertanha, 2020). To the best of our knowledge, our paper is the first to prove RD identification of returns of school assignment in matching mechanisms with strategically reported preferences. We note that our two-step approach differs from the usual control function approach because our first step partially identifies the control variable instead of point-identifying it as in the usual approach. For this reason, we call it control mapping approach. Various other examples in econometrics feature this structure, see Han and Kaido (2025). Our paper unifies and complements two branches of the literature. One branch features the methods proposed by Agarwal and Somaini (2018) and Fack et al. (2019) that take into account strategic reporting and identify students’ true preferences, but they do not focus on causal effects of school matches on future outcomes. The other branch features the methods of Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022) that identify the causal effects of different assignments but control for reported instead of true preferences. An RD strategy that controls for reported preferences identifies causal parameters that could be useful for a class of counterfactuals that differs from the one we consider in this paper. It is important to note, however, that in some empirical contexts, agents may have precise ex ante information about ex post matching cutoffs at the time when they choose preferences to report. In this case, misreporting behavior may arise near a cutoff and this confounds identification in an RD that controls for reported preferences; we find that controlling for true preferences makes the RD strategy robust to this possibility. In the same branch of the literature, Chen (2023) provides a comprehensive set of causal parameters that are identified when assignment is based both on lottery- and RD-driven variation.
@@
We apply our two-step identification strategy to matching data from Chile. Chile has a centralized DA mechanism that assigns students to university–major pairs. In 2010, 88,000 students ranked at most eight university–major pairs out of a total of 1,092 options available. Thus, the mechanism constrains students’ preference rankings, and students have incentives to behave strategically. The methods proposed by Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022) are not directly applicable in the Chilean case, even if all students report their preferences truthfully, as we explain below.
@@
The methods of Kirkeboen et al. (2016) apply to the SD mechanism—a particular case of the DA mechanism in which all schools utilize the same placement score. In the SD case, the counterfactual set of schools for students just above or just below a cutoff does not vary across students. For example, for students just above a cutoff, their counterfactual set includes all schools whose cutoffs are lower than the cutoff in question. The same does not apply in DA, the mechanism used in Chile. Students have multiple placement scores, and the set of feasible schools may vary widely across students. Defining the counterfactual set is an important step in the RD identification strategy because local preferences are defined over these sets, which then become the control variable in the RD strategy.
@@
Abdulkadiroglu et al. (2022) propose a solution to the DA counterfactual problem by constructing a propensity score control variable. They study New York City public high schools, where placement scores are functions of integer priority scores plus a continuously distributed variable with full support. This particular structure of school priorities does not correspond to the Chilean case, where program-specific placement scores are computed as functions of five primitive scores: math, language, history, science, and an average score from high school. These functions are different across programs and sometimes nonlinear. This setting requires the counterfactual set of schools to be carefully defined such that controlling for local preferences does not violate the continuity assumptions required by RD. We therefore propose a general method that applies to such empirical contexts and leads to point identification if students are truthful but partial identification otherwise.
We illustrate our methods using data from college applications in Chile, beginning by documenting clear evidence of strategic behavior among applicants. The analysis proceeds in two parts. First, we estimate average structural functions in the context of the Bachillerato de Ingreso Común at the University of Chile—a selective entry program with a large applicant pool that allows students to explore multiple disciplines before declaring a major, and is often viewed as an alternative route into competitive programs. We examine whether assignment to this program affects two outcomes: graduating from one’s original local first-best program $j$ and graduating from a top university. The results reveal substantial heterogeneity across $j$, indicating that students’ preferences matter for graduation outcomes and are correlated with unobserved factors such as effort or ability not captured by test scores. Second, we estimate treatment effects focusing on Medicine at PUC Santiago, a program whose high selectivity, popularity, and predictable cutoff make strategic behavior especially likely. We measure the effect of being assigned to Medicine rather than to common next-best alternatives $k$ on the two graduation outcomes above. In all cases, our preferred bounds identify a positive effect, with magnitudes that vary across alternatives $k$.
@@
The rest of this paper proceeds as follows. Section 2 lays out the matching model for a continuum population of students and a finite number of schools. Section 3 examines point identification of average treatment effects when students are truth-tellers. Section 4 examines partial identification when students strategically report their preferences, with two subsections: Section 4.1 provides tools for construction of local preference sets that apply to constrained DA mechanisms, and Section 4.2 discusses how to use local preference sets constructed in this or other ways to derive bounds on the average treatment effects. We illustrate our identification approach with the Chilean data in Section 5. The appendix presents all proofs for the paper plus additional results.
@@ 2 Model
We consider a continuum population of students and a set of $J$ schools, $\mathcal{J}:=\{1,\ldots,J\}$, that have capacities $\{q_{1},\ldots,q_{J}\}$ defined in terms of shares of the student population (Azevedo and Leshno, 2016). Denote by $\Omega$ the set of all students in the universe of interest and use $\omega$ to index an individual student type. The student type consists of three objects. First, $Q(\omega)$ denotes the true (strict) preference relation of student $\omega$ over the set of options $\mathcal{J}^{0}:=\mathcal{J}\cup\{0\}$, which includes schools $\mathcal{J}$ and an outside option 0. For example, if $J=2$ and $Q(\omega)=\{1,2,0\}$, then 1 is preferred to 2 (i.e., $1Q(\omega)2$), 1 is preferred to 0 (i.e., $1Q(\omega)0$), and 2 is preferred to 0 (i.e., $2Q(\omega)0$). Let $\mathcal{Q}$ be the set of all strict preference relations over $\mathcal{J}^{0}$ that admit at least one school that is acceptable. A school $j\in\mathcal{J}$ is “acceptable” for student $\omega$ if it is preferred to that student’s outside option, i.e., $jQ(\omega)0$. We define $\bar{Q}$ as the weak preference relation induced by $Q$, i.e., $j\bar{Q}k\Leftrightarrow jQk$ or $j=k$. The second object of the student type is a vector of scores $\bm{\mathrm{R}}(\omega):=(R_{1}(\omega),\ldots,R_{J}(\omega))\in\mathcal{R}\subseteq\mathbb{R}^{J}$, where each school $j$ utilizes $R_{j}$ to rank students for admission. The third and last object, $Y(\omega,d)$, is the potential outcome of student $\omega$ if the student is assigned to option $d\in\mathcal{J}^{0}$. Each student has a potential outcome function $Y(\omega,\cdot)$ that maps from $\mathcal{J}^{0}$ to $\mathcal{Y}\subseteq\mathbb{R}$. We call $\bm{\mathrm{\Gamma}}$ the set of all possible potential outcome functions. The set of all student types is $\Omega:=\mathcal{Q}\times\mathcal{R}\times\bm{\mathrm{\Gamma}}$. In a continuum economy, there is a probability measure $\mathbb{P}$ over $\Omega$ and the Borel $\sigma$-algebra of the product space $\Omega$. We suppress the argument $\omega$ whenever it is unnecessary for ease of notation, e.g., $Y(d)$ vs. $Y(\omega,d)$ and $Q$ vs. $Q(\omega)$.
@@
A “matching” is described by a measurable function $\mu:\Omega\to\mathcal{J}^{0}$ that satisfies two conditions: for every $j\in\mathcal{J}$, (i) the mass of students matched to $j$ is less than or equal to the capacity of school $j$, i.e., $\mathbb{P}\{\omega:\mu(\omega)=j\}\leq q_{j}$; and (ii) the set of students who weakly prefer option $j\in\mathcal{J}^{0}$ over their matching, i.e., $\{\omega:j\bar{Q}(\omega)\mu(\omega)\}$, is an open set.444 Azevedo and Leshno (2016) impose the same condition to rule out multiplicity of stable matchings that differ in a set of types with measure zero. For every student type $\omega$, $\mu(\omega)$ is either the school $j$ to which the student is matched or zero. When $\mu(\omega)=0$, the student is unmatched and takes an outside option. An important definition for this paper is that of stability.
@@
A mechanism $\varphi$ matches students to schools by mapping the students’ scores and submitted preference lists to schools. Student $\omega$ submits a preference list $P(\omega)\subseteq\mathcal{J}$, which is an ordered list of her acceptable schools. For example, for $J=3$, if $Q(\omega)=\{1,2,0,3\}$, then $P(\omega)=\{1,2\}$, as long as the student submits her true list of acceptable schools. The number of schools in $P$, denoted $|P|$, is at least one because everyone participating in the match has at least one acceptable school. As with $\bar{Q}$, we also define $\bar{P}$ as the weak preference relation induced by $P$. A mechanism takes as inputs everyone’s submitted preferences (i.e., a correspondence $P:\Omega\rightrightarrows\mathcal{J}$) and everyone’s scores (i.e., a vector-valued function $\bm{\mathrm{R}}:\Omega\to\mathcal{R}$) and gives rise to a matching function. Formally, $\varphi(P,\bm{\mathrm{R}}):\Omega\to\mathcal{J}^{0}$. We say a mechanism $\varphi$ is strategy proof if, for every student, submitting the true ranking of acceptable schools is a weakly dominant strategy—in other words, any student $\omega$’s misreporting of $P$ never leads to a better option and sometimes leads to a worse option, depending on what other students submit. We say a student is a truth-teller if her $P$ equals her true ranking of acceptable schools. Otherwise, we say she is strategic or not a truth-teller.
@@
For placement scores $\bm{\mathrm{S}}:\Omega\to\mathcal{S}\subseteq\mathbb{R}^{J}$, $\bm{\mathrm{S}}(\omega):=(S_{1}(\omega),\ldots,S_{J}(\omega))$ and admission cutoffs $\bm{\mathrm{c}}\in\mathcal{S}$, $\bm{\mathrm{c}}:=(c_{1},\ldots,c_{J})$, the set of feasible options of a student $\omega$ equals all schools for which her placement scores clear the admission cutoffs plus the outside option: $\{0\}\cup\{j\in\mathcal{J}:S_{j}(\omega)\geq c_{j}\}$; student $\omega$’s best feasible option is the option that ranks first according to $Q(\omega)$ among her feasible options. We say the matching $\mu:\Omega\to\mathcal{J}^{0}$ has cutoff characterization if there exist placement scores $\bm{\mathrm{S}}:\Omega\to\mathcal{S}$ and admission cutoffs $\bm{\mathrm{c}}\in\mathcal{S}$ such that, for every $\omega\in\Omega$, the matching $\mu(\omega)$ equals student $\omega$’s best feasible option according to $Q(\omega)$.